log3(x-1) - log3(x+2) = 2
Answer by edjones(7311)

log3(x-1) - log3(x+2) = 2
3^(log3(x-1))-3^log3(x+2)=3^2
(x-1)/(x+2)=9
x-1=9x+18
-8x=19
x=-19/8
But there can't be a log of a negative number. The equation is false.